OpenStudy (anonymous):

The square root of 5 +the square root of 80 = the square root of 5 + square root of 16 times 5 = the square root + 4 and the square root of 5 = 5 and the square root of 5. I am willing to offer medals if who ever helps me is willing to help with a few other problems tomorrow and also explain the process. I will give you medals for each individual problem if you become a fan and help me out with these Radicals and Complex Numbers. Thanks!!!

OpenStudy (anonymous):

√(5) + √(80) = √(5) + √(16*5) = √(5) + __ √(5) = __ Here is a visual of my problem

jimthompson5910 (jim_thompson5910):

You use the rule \[\Large \sqrt{x*y} = \sqrt{x}*\sqrt{y}\]

jimthompson5910 (jim_thompson5910):

so, \[\Large \sqrt{16*5} = \sqrt{16}*\sqrt{5}\] \[\Large \sqrt{16*5} = ??\]

OpenStudy (anonymous):

I guess? Im super confused I haven't worked on a problem like this in a while. If your familiar with my math lab then that's where I've gotten this problem from.

OpenStudy (anonymous):

if you can explain the process of how you get the answer too then that will be great as well.

jimthompson5910 (jim_thompson5910):

what is the square root of 16 equal to?

OpenStudy (anonymous):

the square root of 4 then the square root of 2

jimthompson5910 (jim_thompson5910):

the square root of 16 is 4 once you take the square root, the square root symbol goes away

jimthompson5910 (jim_thompson5910):

\[\Large \Large \sqrt{16*5} = \sqrt{16}*\sqrt{5}\] \[\Large \Large \sqrt{16*5} = 4\sqrt{5}\]

OpenStudy (anonymous):

THe answer is supposed to be 5 on the square root of 5 im just confused on why that is the answer and how they are getting the answer so that I can other questions like one

OpenStudy (anonymous):

5√(5)

jimthompson5910 (jim_thompson5910):

so do you see how \[\Large \Large \sqrt{16*5} = 4\sqrt{5}\] ??

OpenStudy (anonymous):

you took the square root of each one individually right?

jimthompson5910 (jim_thompson5910):

yes, the square root of 16 is 4 the square root of 5 is some decimal, so we leave that alone

jimthompson5910 (jim_thompson5910):

so, \[\Large \sqrt{5}+\sqrt{16*5} = \sqrt{5}+4\sqrt{5}\]

OpenStudy (anonymous):

yes bc its prime

jimthompson5910 (jim_thompson5910):

there is a 1 in front of that first sqrt(5) term \[\Large \sqrt{5}+4\sqrt{5} = 1\sqrt{5}+4\sqrt{5}\]

jimthompson5910 (jim_thompson5910):

the last step is to add the numbers outside the root 5 terms 1+4 = 5 think of it as 1x+4x = 5x where x is equal to sqrt(5)

OpenStudy (anonymous):

oh okay i think i get it.

jimthompson5910 (jim_thompson5910):

so you see how \[\Large 1\sqrt{5}+4\sqrt{5} = 5\sqrt{5}\] ?

OpenStudy (anonymous):

yes I got the correct answer on another problem! thank you!! will you be on tomorrow?

jimthompson5910 (jim_thompson5910):

yes I'll be online tomorrow

jimthompson5910 (jim_thompson5910):

and I'm glad to be of help and that it's making more sense now

OpenStudy (anonymous):

if so i will give you several other medals for each problem that i have

OpenStudy (anonymous):

perfect!! okay tomorrow i will put on a few other problems for us to work with see ya tomorrow jim_thompson5910

jimthompson5910 (jim_thompson5910):

ok sounds good, have a good day/night